# Autonomous Driving in Unstructured Environments: How Far Have We Come?

### v3

@@ IV-B2 Sample-based Methods
Jaillet et al. [30] T-RRT
Jiang et al. [31] R2-RRT*
Yin et al. [209] ER-RRT*, Surrogate Modeling
Tian et al. [210] Potential Field-based RRT*
Zheng et al. [211] PRM, A*
The PRM algorithm first randomly selects nodes in the search space, then connects all nodes, including the start and end points, and performs collision detection to form a collision-free connected graph. Finally, it uses a search algorithm to find the shortest path connecting the start and end points. The PRM algorithm also has probabilistic completeness and avoids a complete search of the entire area through sampling, thereby reducing the computational load. Thus, this algorithm is more efficient and simpler to implement than grid search methods, and it is widely used in high-dimensional space planning problems. Zheng et al. [211] propose a two-stage path planning algorithm for complex off-road environments. They used an improved PRM algorithm to quickly and roughly determine the initial path, and the A* algorithm for precise path planning.
The RRT algorithm is a sampling-based path planning method, especially suitable for path search problems in high-dimensional spaces. Its core idea is to quickly expand a tree structure through random sampling to find a feasible path from the starting point to the target point. The RRT algorithm is simple to operate, has fast solution speed, and good real-time performance, and it is widely used in high-dimensional spaces or large-scale maps. Jaillet et al. [30] propose T-RRT, which can generate not only feasible but also high-quality paths. T-RRT considers a user-given cost function defined over the configuration space as an additional input to the standard path-planning problem. For example, in outdoor navigation problems, the cost map may correspond to the elevation map of the terrain to compute movements that minimize climbs in steep slope areas.
@@
Through path reconnection and optimization mechanisms, the RRT* algorithm can find near-optimal paths in the search space. Tian et al. [210] propose a potential field-based RRT* motion planning algorithm for vehicles to avoid risks, constructing a configuration space with potential fields to determine risk ranges around obstacles and off-road terrain. Due to its random sampling characteristics, the RRT* algorithm has global search capabilities, avoiding local optima. However, the optimization process requires additional computation time, and the complexity of nearest node and path feasibility detection may result in less ideal performance in applications with high real-time requirements. Jiang et al. [31] propose state mobility reliability and task mobility reliability to quantify the mobility reliability of AGVs. They developed two reliability-based robust mission planning models and generated optimal paths that meet specific reliability requirements. Yin et al. [209] combine surrogate modeling with the RRT* algorithm and proposed a novel and efficient reliability-based global path planning method.
IV-B3 Biologically-inspired Methods
@@
TABLE XII: Biologically-inspired path planning methods for unstructured environments.
@@
Wang et al. [32] ACO
Zhu et al. [33] ACO, APF
Peng et al. [216] GA, PSO
Hu and Yang [217] Knowledge-based GA
Li et al. [218] Improved GA
Ant Colony Optimization: Simulates the foraging behavior of ants, where ants leave pheromone trails to mark paths and use positive feedback to find the optimal path. As ants move, they deposit pheromones, which evaporate over time; paths with higher pheromone concentrations attract more ants, gradually forming the optimal path. Table XII details the biologically-inspired path planning methods for unstructured environments. Wang et al. [32] propose an improved ant colony path planning algorithm after analyzing the combined impact of terrain slope and soil strength on vehicle off-road passability. Zhu et al. [33] combine the artificial potential field method with ACO, introducing inducible heuristic factors and dynamically adjusting the state transition rules of ACO, resulting in higher global search capabilities and faster convergence.
Particle Swarm Optimization: Simulates the behavior of bird flocks, finding the optimal solution through cooperation and competition among particles. Each particle represents a candidate solution, adjusting its position based on its own experience and the experience of the swarm, gradually converging to the optimal solution. Peng et al. [216] represent vehicle ride comfort, road-holding, and handling performance in rugged terrain through the weighted root mean square value of sprung mass vertical acceleration, suspension travel, and tire deformation, optimizing these objectives using GA, PSO, and a GA-PSO hybrid algorithm.
Genetic Algorithm: Simulates natural selection and genetic variation processes, evolving to the optimal solution through iterative selection, recombination, and mutation of individuals in the population. Individuals in the population represent candidate solutions, with higher fitness individuals more likely to reproduce and pass on their characteristics to the next generation. Hu and Yang [217] propose a novel knowledge-based genetic algorithm for robot path planning in complex unstructured environments, developing multiple operators to enhance the algorithm’s performance, such as combining local path planning operators. Li et al. [218] propose an improved genetic algorithm to address the shortcomings of basic genetic algorithms.
@@
IV-C Local Path Planning
IV-C1 Optimization-based methods
Optimization-based methods transform path planning problems into mathematical optimization problems by defining an objective function (such as path length, energy consumption, time, etc.) and constraints (such as avoiding obstacles, path smoothness, etc.), and then use optimization algorithms to find the optimal path. In many cases, optimization methods can find the global optimum or a solution close to it. The objective function and constraints can be adjusted according to specific problem requirements, providing great flexibility. However, the computational complexity of solving complex optimization problems can be high, especially for nonlinear and mixed integer optimization problems. Common optimization algorithms include Linear Programming (LP)[219], Nonlinear Programming (NLP)[220], and Mixed Integer Programming (MIP)[221], as shown in Table XIII.
TABLE XIII: Optimization-based methods for unstructured environments.
@@ IV-C1 Optimization-based methods
Lee and Wah [34] SNOPT,FilMINT, MINLP BB
Guo et al. [35] NLP, CPLEX, A*
Zang et al. [222] MIQP, BFS, STMC
@@
Lee and Wah [34] develop heuristics for finding good starting points in solving large-scale nonlinear constrained optimization problems, especially NLP and MINLP with nonlinear non-convex functions. The experimental evaluations on NLP and MINLP benchmark problems show that their approach can solve more problems with fewer iterations compared to the best existing solvers from their default starting points.
Guo et al. [35] propose a down-sized initialization strategy for optimization-based unstructured trajectory planning, which only optimizes critical variables to provide a proper initial guess for the trajectory planning problem, and they demonstrate that this strategy can facilitate the solution process of the planner and improve the solution time and trajectory quality through experiments.
Zang et al. [222] propose a coordinated behavior planning and trajectory planning framework for multiple unmanned ground vehicles in unstructured narrow interaction scenarios, which integrates an efficient behavior planning layer based on MIQP and BFS and an optimal trajectory planning layer with a spatio-temporal motion corridor approach, and they validate the framework in simulations and real-vehicle experiments, demonstrating that it can generate safe, smooth, and efficient trajectories compared to existing planning methods.
IV-C2 Artificial Potential Field
@@
TABLE XIV: Artificial potential field-based path planning for unstructured environments.
@@
Yao et al. [36] Improved Black Hole Potential Fields, RL
Jiang et al. [37] APF, Coarse2fine A*
Linghong et al. [225] APF
@@
The artificial potential field method is relatively simple to implement, has low computational overhead, and thus is efficient. Additionally, since the attractive and repulsive fields are manually designed, it offers strong scalability, allowing multiple factors to be considered during planning. Table XIV lists the artificial potential field-based path planning methods for unstructured environments. Jiang et al. [37] adopt the concept of artificial potential fields to characterize various risks and incorporate the risk assessment results into the planning module of unmanned ground vehicles. Based on this, they proposed a global planning algorithm called Coarse2fine A*, which considers risks and improves efficiency and flexibility.
Yao et al. [36] combine improved black hole potential fields with reinforcement learning, where the black hole potential field is used as the environment in the reinforcement learning algorithm. The agent automatically adapts to the environment and learns how to use basic environmental information to find the target. However, the main limitation of this method is its tendency to get trapped in local optima. This means that during the potential energy descent search process, the algorithm may stall at a local minimum and fail to reach the goal. Furthermore, if an obstacle is near the goal, its repulsive field may interfere with the potential energy of the goal, causing the goal not to be a global minimum and leading to incorrect goal identification. Linghong et al. [225] construct driving safety risks in complex and variable urban tunnels based on the theory of artificial potential fields.
IV-C3 Dynamic Window Approach-based Methods
TABLE XV: Dynamic window approach-based path planning methods for unstructured environments.
@@ IV-C3 Dynamic Window Approach-based Methods
Yang et al. [38] DWA, Improved A*
Eirale et al. [39] DWA, DRL
Zhang and Li [226] DWA, Improved D* Lite Algorithm
@@
The DWA algorithm can quickly compute the optimal path in dynamically changing environments. The obstacle distance metric in the evaluation function effectively prevents collisions, ensuring the safety of the path planning. Table XV details the path planning methods based on the dynamic window approach for unstructured environments. Yang et al. [38] propose an improved A* algorithm and DWA to enhance the accuracy and speed of tracked vehicle automatic parking. However, due to its local search nature, the DWA algorithm may fall into local optima and thus cannot guarantee finding the globally optimal path. Eirale et al. [39] also combine deep reinforcement learning with DWA, introducing a new person-following method where the robot’s linear velocity was calculated by DWA, and the angular velocity was predicted by the deep reinforcement learning algorithm. Zhang and Li [226] use an improved D* Lite algorithm to generate global waypoints for path planning and employed DWA for local path planning, where DWA parameters were selected by a deep reinforcement learning algorithm.
IV-C4 Data-driven Methods
@@
TABLE XVI: Data-driven path planning methods for unstructured environments.
@@
Kaushik and Madhava Krishna [227] DDPG
Liu et al. [40] Evolutionary RL
Semnani et al. [228] GA3C-CADRL, FMP
Ahn et al. [229] Imitation Learning
Wang et al. [41] PPO, Curriculum Learning
Li et al. [230] Deep Learning, Quadratic Optimization
Ginerica et al. [231] RNN, DWA
@@
Learning-based path planning in unstructured environments utilizes neural networks’ efficient learning and fitting capabilities in complex environments. Deep neural network models extract features from sensor data (e.g., Cameras, LiDAR) and learn effective path planning strategies [232]. For instance, CNNs [233] can extract terrain features, recognize obstacles, and identify roads, aiding the robot in determining the optimal path. Deep learning models handle complex unstructured environments, such as mountains and forests, by learning path planning directly from sensor data without the need for manual feature engineering. They can be adjusted and optimized according to different environments and tasks. However, deep learning models typically require a large amount of labeled data for training, especially in complex environments. Training and inference of deep learning models demand substantial computational resources, posing challenges in real-time applications. Moreover, their generalization ability in unseen environments or extreme conditions may be limited. Table XVI lists the data-driven path planning methods for unstructured environments.
@@
Ahn et al. [229] propose a method for autonomous driving in unstructured environments using vision-based occupancy grid maps, achieved through imitation learning with expert driving data to train deep neural networks. Li et al. [230] demonstrate the potential of a trajectory planning method based on deep learning and optimization in the practical application of autonomous driving in mining areas. Ginerica et al. [231] also propose a vision dynamics approach for path planning and navigation of quadruped robots in unstructured environments, specifically forest roads, using a recurrent neural network that employs RGB-D sensors to construct a sequence of previous depth sensor observations and predict future observations over a limited time span. Xu et al. [234] propose an end-to-end transformer networks based approach for map-less autonomous driving. The proposed model takes raw LiDAR data and noisy topometric map as input and produces precise local trajectory for navigation. It demonstrates the effectiveness in real-world driving data, including both urban and rural areas.
@@
Deep reinforcement learning (DRL) [235] in path planning combines deep learning and reinforcement learning techniques to enable agents to autonomously learn optimal paths in complex environments. Specifically, DRL uses neural networks to handle high-dimensional state spaces. The input to the neural network includes various environmental information, such as obstacles and target locations, while the output is the agent’s action strategy. The core of reinforcement learning is to obtain feedback through interactions with the environment, i.e., rewards and penalties, to continuously update and optimize the strategy. Kaushik and Madhava Krishna [227] propose using the DDPG [236] algorithm to enable vehicles to learn various driving behaviors in unstructured environments through simulation. Agents improve their behavior through trial-and-error learning, maximizing total rewards by accumulating experience. In path planning applications, DRL effectively handles dynamic, unknown, and complex environments. Agents do not need to pre-know all environmental information; they can find the optimal path from start to target through continuous exploration and learning.
@@
Liu et al. [40] propose MAPPER, a decentralized partially observable multi-agent path planning method based on evolutionary reinforcement learning, to learn effective local planning strategies in mixed dynamic environments. Semnani et al. [228] develop a hybrid algorithm combining deep reinforcement learning and Force-based Motion Planning (FMP) to address distributed motion planning in dense dynamic environments. Their hybrid approach employs the simpler FMP method in stuck, simple, and high-risk situations, and continues to use RL in normal circumstances where FMP cannot generate optimal paths. Wang et al. [41] introduce a path planning method for unstructured environments based on deep reinforcement learning, where agents are trained in a low-dimensional simulator constructed from occupancy grid maps. A self-supervised pre-text task helps agents learn environmental knowledge, and an adaptive curriculum learning method improves training efficiency and generalization ability.
V Motion Control

### v4

@@ 4.2.2 Sample-based Methods
[75] T-RRT
[76] R2-RRT*
[238] ER-RRT*, Surrogate Modeling
[186] Potential Field-based RRT*
[251] PRM, A*
The PRM algorithm first randomly selects nodes in the search space, then connects all nodes, including the start and end points, and performs collision detection to form a collision-free connected graph. Finally, it uses a search algorithm to find the shortest path connecting the start and end points. The PRM algorithm also has probabilistic completeness and avoids a complete search of the entire area through sampling, thereby reducing the computational load. Thus, this algorithm is more efficient and simpler to implement than grid search methods, and it is widely used in high-dimensional space planning problems. [251] propose a two-stage path planning algorithm for complex off-road environments. They used an improved PRM algorithm to quickly and roughly determine the initial path, and the A* algorithm for precise path planning.
The RRT algorithm is a sampling-based path planning method, especially suitable for path search problems in high-dimensional spaces. Its core idea is to quickly expand a tree structure through random sampling to find a feasible path from the starting point to the target point. The RRT algorithm is simple to operate, has fast solution speed, and good real-time performance, and it is widely used in high-dimensional spaces or large-scale maps. [75] propose T-RRT, which can generate not only feasible but also high-quality paths. T-RRT considers a user-given cost function defined over the configuration space as an additional input to the standard path-planning problem. For example, in outdoor navigation problems, the cost map may correspond to the elevation map of the terrain to compute movements that minimize climbs in steep slope areas.
@@
Through path reconnection and optimization mechanisms, the RRT* algorithm can find near-optimal paths in the search space. [186] propose a potential field-based RRT* motion planning algorithm for vehicles to avoid risks, constructing a configuration space with potential fields to determine risk ranges around obstacles and off-road terrain. Due to its random sampling characteristics, the RRT* algorithm has global search capabilities, avoiding local optima. However, the optimization process requires additional computation time, and the complexity of nearest node and path feasibility detection may result in less ideal performance in applications with high real-time requirements. [76] propose state mobility reliability and task mobility reliability to quantify the mobility reliability of AGVs. They developed two reliability-based robust mission planning models and generated optimal paths that meet specific reliability requirements. [238] combine surrogate modeling with the RRT* algorithm and proposed a novel and efficient reliability-based global path planning method.
4.2.3 Biologically-inspired Methods
@@
Table 12: Biologically-inspired path planning methods for unstructured environments.
@@
[199] ACO
[255] ACO, APF
[150] GA, PSO
[71] Knowledge-based GA
[102] Improved GA
Ant Colony Optimization: Simulates the foraging behavior of ants, where ants leave pheromone trails to mark paths and use positive feedback to find the optimal path. As ants move, they deposit pheromones, which evaporate over time; paths with higher pheromone concentrations attract more ants, gradually forming the optimal path. Table 12 details the biologically-inspired path planning methods for unstructured environments. [199] propose an improved ant colony path planning algorithm after analyzing the combined impact of terrain slope and soil strength on vehicle off-road passability. [255] combine the artificial potential field method with ACO, introducing inducible heuristic factors and dynamically adjusting the state transition rules of ACO, resulting in higher global search capabilities and faster convergence.
Particle Swarm Optimization: Simulates the behavior of bird flocks, finding the optimal solution through cooperation and competition among particles. Each particle represents a candidate solution, adjusting its position based on its own experience and the experience of the swarm, gradually converging to the optimal solution. [150] represent vehicle ride comfort, road-holding, and handling performance in rugged terrain through the weighted root mean square value of sprung mass vertical acceleration, suspension travel, and tire deformation, optimizing these objectives using GA, PSO, and a GA-PSO hybrid algorithm.
Genetic Algorithm: Simulates natural selection and genetic variation processes, evolving to the optimal solution through iterative selection, recombination, and mutation of individuals in the population. Individuals in the population represent candidate solutions, with higher fitness individuals more likely to reproduce and pass on their characteristics to the next generation. [71] propose a novel knowledge-based genetic algorithm for robot path planning in complex unstructured environments, developing multiple operators to enhance the algorithm’s performance, such as combining local path planning operators. [102] propose an improved genetic algorithm to address the shortcomings of basic genetic algorithms.
@@
4.3 Local Path Planning
4.3.1 Optimization-based methods
Optimization-based methods transform path planning problems into mathematical optimization problems by defining an objective function (such as path length, energy consumption, time, etc.) and constraints (such as avoiding obstacles, path smoothness, etc.), and then use optimization algorithms to find the optimal path. In many cases, optimization methods can find the global optimum or a solution close to it. The objective function and constraints can be adjusted according to specific problem requirements, providing great flexibility. However, the computational complexity of solving complex optimization problems can be high, especially for nonlinear and mixed integer optimization problems. Common optimization algorithms include Linear Programming (LP)[26], Nonlinear Programming (NLP)[9], and Mixed Integer Programming (MIP)[1], as shown in Table 13.
Table 13: Optimization-based methods for unstructured environments.
@@ 4.3.1 Optimization-based methods
[99] SNOPT,FilMINT, MINLP BB
[60] NLP, CPLEX, A*
[242] MIQP, BFS, STMC
@@
[99] develop heuristics for finding good starting points in solving large-scale nonlinear constrained optimization problems, especially NLP and MINLP with nonlinear non-convex functions. The experimental evaluations on NLP and MINLP benchmark problems show that their approach can solve more problems with fewer iterations compared to the best existing solvers from their default starting points.
[60] propose a down-sized initialization strategy for optimization-based unstructured trajectory planning, which only optimizes critical variables to provide a proper initial guess for the trajectory planning problem, and they demonstrate that this strategy can facilitate the solution process of the planner and improve the solution time and trajectory quality through experiments.
[242] propose a coordinated behavior planning and trajectory planning framework for multiple unmanned ground vehicles in unstructured narrow interaction scenarios, which integrates an efficient behavior planning layer based on MIQP and BFS and an optimal trajectory planning layer with a spatio-temporal motion corridor approach, and they validate the framework in simulations and real-vehicle experiments, demonstrating that it can generate safe, smooth, and efficient trajectories compared to existing planning methods.
4.3.2 Artificial Potential Field
@@
Table 14: Artificial potential field-based path planning for unstructured environments.
@@
[236] Improved Black Hole Potential Fields, RL
[77] APF, Coarse2fine A*
[113] APF
@@
The artificial potential field method is relatively simple to implement, has low computational overhead, and thus is efficient. Additionally, since the attractive and repulsive fields are manually designed, it offers strong scalability, allowing multiple factors to be considered during planning. Table 14 lists the artificial potential field-based path planning methods for unstructured environments. [77] adopt the concept of artificial potential fields to characterize various risks and incorporate the risk assessment results into the planning module of unmanned ground vehicles. Based on this, they proposed a global planning algorithm called Coarse2fine A*, which considers risks and improves efficiency and flexibility.
[236] combine improved black hole potential fields with reinforcement learning, where the black hole potential field is used as the environment in the reinforcement learning algorithm. The agent automatically adapts to the environment and learns how to use basic environmental information to find the target. However, the main limitation of this method is its tendency to get trapped in local optima. This means that during the potential energy descent search process, the algorithm may stall at a local minimum and fail to reach the goal. Furthermore, if an obstacle is near the goal, its repulsive field may interfere with the potential energy of the goal, causing the goal not to be a global minimum and leading to incorrect goal identification. [113] construct driving safety risks in complex and variable urban tunnels based on the theory of artificial potential fields.
4.3.3 Dynamic Window Approach-based Methods
Table 15: Dynamic window approach-based path planning methods for unstructured environments.
@@ 4.3.3 Dynamic Window Approach-based Methods
[235] DWA, Improved A*
[35] DWA, DRL
[248] DWA, Improved D* Lite Algorithm
@@
The DWA algorithm can quickly compute the optimal path in dynamically changing environments. The obstacle distance metric in the evaluation function effectively prevents collisions, ensuring the safety of the path planning. Table 15 details the path planning methods based on the dynamic window approach for unstructured environments. [235] propose an improved A* algorithm and DWA to enhance the accuracy and speed of tracked vehicle automatic parking. However, due to its local search nature, the DWA algorithm may fall into local optima and thus cannot guarantee finding the globally optimal path. [35] also combine deep reinforcement learning with DWA, introducing a new person-following method where the robot’s linear velocity was calculated by DWA, and the angular velocity was predicted by the deep reinforcement learning algorithm. [248] use an improved D* Lite algorithm to generate global waypoints for path planning and employed DWA for local path planning, where DWA parameters were selected by a deep reinforcement learning algorithm.
4.3.4 Data-driven Methods
@@
Table 16: Data-driven path planning methods for unstructured environments.
@@
[82] DDPG
[118] Evolutionary RL
[165] GA3C-CADRL, FMP
[2] Imitation Learning
[204] PPO, Curriculum Learning
[104] Deep Learning, Quadratic Optimization
[52] RNN, DWA
@@
Learning-based path planning in unstructured environments utilizes neural networks’ efficient learning and fitting capabilities in complex environments. Deep neural network models extract features from sensor data (e.g., Cameras, LiDAR) and learn effective path planning strategies [95]. For instance, CNNs [54] can extract terrain features, recognize obstacles, and identify roads, aiding the robot in determining the optimal path. Deep learning models handle complex unstructured environments, such as mountains and forests, by learning path planning directly from sensor data without the need for manual feature engineering. They can be adjusted and optimized according to different environments and tasks. However, deep learning models typically require a large amount of labeled data for training, especially in complex environments. Training and inference of deep learning models demand substantial computational resources, posing challenges in real-time applications. Moreover, their generalization ability in unseen environments or extreme conditions may be limited. Table 16 lists the data-driven path planning methods for unstructured environments.
@@
[2] propose a method for autonomous driving in unstructured environments using vision-based occupancy grid maps, achieved through imitation learning with expert driving data to train deep neural networks. [104] demonstrate the potential of a trajectory planning method based on deep learning and optimization in the practical application of autonomous driving in mining areas. [52] also propose a vision dynamics approach for path planning and navigation of quadruped robots in unstructured environments, specifically forest roads, using a recurrent neural network that employs RGB-D sensors to construct a sequence of previous depth sensor observations and predict future observations over a limited time span. [226] propose an end-to-end transformer networks based approach for map-less autonomous driving. The proposed model takes raw LiDAR data and noisy topometric map as input and produces precise local trajectory for navigation. It demonstrates the effectiveness in real-world driving data, including both urban and rural areas.
@@
Deep reinforcement learning (DRL) [205] in path planning combines deep learning and reinforcement learning techniques to enable agents to autonomously learn optimal paths in complex environments. Specifically, DRL uses neural networks to handle high-dimensional state spaces. The input to the neural network includes various environmental information, such as obstacles and target locations, while the output is the agent’s action strategy. The core of reinforcement learning is to obtain feedback through interactions with the environment, i.e., rewards and penalties, to continuously update and optimize the strategy. [82] propose using the DDPG [110] algorithm to enable vehicles to learn various driving behaviors in unstructured environments through simulation. Agents improve their behavior through trial-and-error learning, maximizing total rewards by accumulating experience. In path planning applications, DRL effectively handles dynamic, unknown, and complex environments. Agents do not need to pre-know all environmental information; they can find the optimal path from start to target through continuous exploration and learning.
@@
[118] propose MAPPER, a decentralized partially observable multi-agent path planning method based on evolutionary reinforcement learning, to learn effective local planning strategies in mixed dynamic environments. [165] develop a hybrid algorithm combining deep reinforcement learning and Force-based Motion Planning (FMP) to address distributed motion planning in dense dynamic environments. Their hybrid approach employs the simpler FMP method in stuck, simple, and high-risk situations, and continues to use RL in normal circumstances where FMP cannot generate optimal paths. [204] introduce a path planning method for unstructured environments based on deep reinforcement learning, where agents are trained in a low-dimensional simulator constructed from occupancy grid maps. A self-supervised pre-text task helps agents learn environmental knowledge, and an adaptive curriculum learning method improves training efficiency and generalization ability.
Errors in traversability estimation, such as misclassifying soft soil or mud as drivable, can propagate to planning modules and increase the risk of collisions or vehicle immobilization. Future work should explore uncertainty-aware planning and other mitigation strategies to account for perception inaccuracies, enhancing the overall robustness of off-road autonomous navigation.
5 Motion Control